A farmer has a rectangular garden plot surrounded by 200 ft of fence. Find the lenght and with of the gfarden if its area is 2400ft(square)
Answer by Nate(3500)

2w + 2l = perimeter
2w = 200 - 2l
w = 100 - l
lw = area
(100 - l)(l) = 2400
-l^2 + 100l - 2400 = 0
l^2 - 100l + 2400 = 0
(l - 40)(l - 60) = 0
length = 40 or 60
plug
w x 40
w = 100 - l
w = 100 - 40
w = 60
60 x 40 ? width is shorter than length (this is a not true answer)
True Answer
w x 60
w = 100 - l
w = 100 - 60
w = 40
40 x 60